Alpha Inc. is evaluating the purchase of a machine costing $350,000. The expected useful life of the machine is 5 years at the end which it would have no residual value, and the depreciation is assumed to be on straight-line basis. The estimated total income from the machine is $500,000. The expected average rate of return for this machine is:
Answer #1

Average net income = 500000/5 = 100000

Average investment = 350000/2 = 175000

Average rate of return = Average net income/Average investment

= 100000/175000

Average rate of return = 57.14%

Note : Sometime it can be = 100000/350000 = 28.57% but it could prefer very less